# P3-LLM: An Integrated NPU-PIM Accelerator for LLM Inference Using Hybrid Numerical Formats

### v1

@@ VI-C Hardware Performance
Comparison with Existing Low-Precision PIM. We compare the latency of P3-LLM with Pimba [pimba], a SoTA low-precision PIM architecture adopting the 8-bit microscaling format [microscaling]. The original Pimba employs KV-cache-only quantization since it targets cloud serving scenarios, where KV-cache dominates the memory footprint. Hence, we also examine an enhanced version of Pimba with 8-bit weight-activation quantization. To ensure fairness, we match the timing parameters and the number of HBM stacks across all systems. Fig. 11 presents the normalized decoding performance of P3-LLM and Pimba under batch sizes of 2 and 4. The original Pimba has the lowest performance because at small batch sizes, weights can dominate the overall memory footprint but remain unquantized. By adopting 8-bit weight-activation quantization, on average, the enhanced Pimba achieves $2.1\times$ better performance compared to its original design. The proposed P3-LLM further yields an average of $3.4\times$ performance boost compared to the enhanced Pimba. These speedups are attributed to the more aggressive W4A8KV4P8 quantization and the high-throughput PCU design. By quantizing weights and KV-cache to 4 bits with minimal accuracy loss, P3-LLM reduces the memory access compared to Pimba. Moreover, the throughput-enhanced PCU of P3-LLM enables temporal input reuse to double the computational throughput, facilitating efficient execution of low-batch linear layers and GQA.

### v2

@@ VI-C Hardware Performance
Comparison with Existing Low-Precision PIM. We compare the performance of P3-LLM with Pimba [pimba], a SoTA low-precision PIM architecture adopting the 8-bit microscaling format [microscaling]. The original Pimba employs KV-cache-only quantization since it targets cloud serving scenarios, where KV-cache dominates the memory footprint. Hence, we also examine an enhanced version of Pimba with 8-bit weight-activation quantization. Fig. 11 presents the normalized decoding speed of P3-LLM and Pimba under batch sizes of 2 and 4. The original Pimba has the lowest performance because at small batch sizes, weights can dominate the overall memory footprint but remain unquantized. By adopting 8-bit weight-activation quantization, on average, the enhanced Pimba achieves $2.1\times$ better performance compared to its original design. The proposed P3-LLM further yields an average of $3.4\times$ performance boost compared to the enhanced Pimba. By quantizing weights and KV-cache to 4 bits with minimal accuracy loss, P3-LLM reduces the memory access compared to Pimba. Moreover, the throughput-enhanced PCU of P3-LLM enables temporal input reuse to double the computational throughput, facilitating efficient execution of low-batch linear layers and GQA.